OpenStudy (jamaljunior):

Mrs. Reed is decorating wreaths. She needs 24 in. of gold ribbon and 45 in. of white ribbon to tie around each wreath. The white ribbon is sold by the foot. About how many feet of white ribbon will Mrs. Reed need to make 18 wreaths?

OpenStudy (awesomedude2078):

what are your choices

OpenStudy (jamaljunior):

its a fill in

OpenStudy (awesomedude2078):

i got 67.5 so maybe you should round it to 68

OpenStudy (jamaljunior):

idk it doesnt say round. ill wait for someones confirmation tho

OpenStudy (awesomedude2078):

ok

563blackghost (563blackghost):

You are told that you require `45 in.` of white ribbon to make one wreath. Now you want to know how much in `FEET` to make 18 wreaths....So you would first multiply `45 in by 18`.... \(\huge{45 \times 18 = ?}\)

563blackghost (563blackghost):

Im sorry I can't confirm if the answer is right since no explanation was made by @awesomedude2078 so I have to right an explanation to make sure you understand how to solve....

563blackghost (563blackghost):

write*

OpenStudy (jamaljunior):

what do i do next?

563blackghost (563blackghost):

What did you find as your answer?

OpenStudy (jamaljunior):

i dont have 1

563blackghost (563blackghost):

Simplify `45 x 18`.

OpenStudy (jamaljunior):

810?

563blackghost (563blackghost):

Correct :) Now we have the `TOTAL INCHES` but we need to find the `TOTAL FEET` So we would divide that by how many inches are in a foot so that would be 12 inches.... So we would divide 810 by 12.... \(\huge{\frac{810}{12}=?}\)

OpenStudy (francesca_marie):

24x18 = 432 gold in total 45x18 = 810 white in total 12 inches = 1 foot 432/12 = 36 feet gold 810/12 = 67.5 feet white - round up to 70 (C). (The question says "about", so it allows for a little flexibility in range.)

OpenStudy (jamaljunior):

oh okay.

OpenStudy (jamaljunior):

thanks @Francesca_Marie and @563blackghost

563blackghost (563blackghost):

np :)

OpenStudy (francesca_marie):

No problem